OpenStudy (anonymous):

Which of the following describes the correct process for solving the equation 3x + 6 = 21 and arrives at the correct solution? Divide both sides by 6 and then add 3. The solution is x = 5. Add 6 to both sides of the equation and then divide by 3. The solution is x = 9. Subtract 6 from both sides of the equation and then divide by 3. The solution is x = 8. Subtract 6 from both sides of the equation and then divide by 3. The solution is x = 5.

OpenStudy (azureilai):

last one

OpenStudy (anonymous):

A water park charges $5 for entry into the park and an additional $2 for each of the big waterslides. Steven spent $17 on his visit to the water park. Which equation can be used to find the number of big waterslides, x, that Steven went on? 5 + 2x = 17 5x + 2 = 17 5x + 2x = 17 5 + 2 = 17x

OpenStudy (anonymous):

Solve y over negative 2 + 5 = 13. - 16 16 -36 36 please help me in this one

OpenStudy (azureilai):

So in the first one, 5 is the constant and the two dollars is dependent on how many slides he take. Which one do you think is the answer?